Consider an example of a new car sale yard and the managers predicted that they would sell 1,400 of model A and 1,800 of model B in the fiscal year. They would prepare their budget accordingly. Sales of model A did not meet expectations and sold 1,200 units. Model B had sales rise to 2,400 vehicles for the year. The expected margin on each vehicle is $2,000 for model A and $3,000 for model B.

Calculate the predicted and actual sales mix and the impact of the variations on income.
1
Expert's answer
2018-08-27T09:29:08-0400
Predicted: 1,400 of product A and 1,800 of product B.
Actual: 1,200 for product A and 2,400 for product B.
The wholesale margin on product A is calculated to be $2,000 and on product B it is $3,000.
The predicted sales mix is 1,400/3,200 = 43.75% of A to 56.25% of B.
The actual sales mix is 1,200/3,600 = 33.33% of A to 66.67% of B.
The sales mix variance for A is 1,200*(0.333 - 0.4375)*$2,000 = -250,800 or unfavorable variance.
For B, the sales mix variance is 2,400*(0.6667 - 0.5625)*$3,000 = 750,240 or favorable variance.